A project is expected to generate the following cash flows: Year Project after-tax cash flows -$350 150 -25 300 The project's cost of capital is 10%, calculate this project’s MIRR.
Answer #1

MIRR = ( FVc / PVfc )1/n -1

where, FVc is the  future value of positive cash flows

and PVfc is the present value of negative cash flows

Cost of Capital = r = 10%

CF0 = -350
CF1 = 150
CF2 = -25
CF3 = 300

PVfc = CF0 + CF2/(1+r)2 = 350 + 25/(1+0.10)2 = $370.66

FVc = CF1(1+r)2 + CF3 = 150(1+0.10)2 + 300 = $481.5

Hence, MIRR = (481.5/370.66)1/3 - 1 = 0.0911 = 9.11%